# CG-HOI: Contact-Guided 3D Human-Object Interaction Generation

### v1

@@ Abstract
We propose CG-HOI, the first method to address the task of generating dynamic 3D human-object interactions (HOIs) from text. We model the motion of both human and object in an interdependent fashion, as semantically rich human motion rarely happens in isolation without any interactions. Our key insight is that explicitly modeling contact between the human body surface and object geometry can be used as strong proxy guidance, both during training and inference. Using this guidance to bridge human and object motion enables generating more realistic and physically plausible interaction sequences, where the human body and corresponding object move in a coherent manner. Our method first learns to model human motion, object motion, and contact in a joint diffusion process, inter-correlated through cross-attention. We then leverage this learned contact for guidance during inference synthesis of realistic, coherent HOIs. Extensive evaluation shows that our joint contact-based human-object interaction approach generates realistic and physically plausible sequences, and we show two applications highlighting the capabilities of our method. Conditioned on a given object trajectory, we can generate the corresponding human motion without re-training, demonstrating strong human-object interdependency learning. Our approach is also flexible, and can be applied to static real-world 3D scene scans.
@@ 1 Introduction
We present CG-HOI, the first approach to address the task of generating realistic 3D human-object interactions from text descriptions, by jointly predicting a sequence of 3D human body motion along with the object motion. Key to our approach is to not only model human and object motion, but also to explicitly model contact as a bridge between human and object. In particular, we model contact by predicting contact distances from the human body surface to the closest point on the surface of the object being interacted with. This explicit modeling of contact helps to encourage human and object motion to be semantically coherent, as well as provide a constraint indicating physical plausibility (e.g., discouraging objects to float without support).
@@ 2 Related Work
Figure 2: Method Overview. Given a brief text description an an object geometry, CG-HOI produces a human-object interaction sequence where both human and object motion are modeled. To produce realistic human-object interactions, we additionally model contact to bridge the interdependent motions. Our method jointly generates all three during training (left), using a U-Net-based diffusion with cross-attention across human, object, and contact. During inference (right), we drive synthesis under guidance of estimated contact to sample more physically plausible interactions.
3D Human Motion Generation.
Generating sequences of 3D humans in motion is a task which evolved noticeably over the last few years. Traditionally, many methods used recurrent approaches [20, 2, 14, 22, 32, 50] and, improving both fidelity and predicted sequence length, graph- and attention-based frameworks [45, 46, 68]. Notably, generation can either happen deterministically, predicting one likely future human pose sequence [20, 50, 45, 46, 18], or stochastically, thereby also modelling the uncertainty inherent to future human motion [4, 7, 47, 86, 91, 17, 9, 85].
@@
3D Human Motion in Scenes.
As human motion typically occurs not in isolation but in the context of an object or surrounding environment, various methods have explored learning plausible placement of humans into scenes, both physically and semantically, [96, 26, 100, 28], forecasting future motion given context [11, 48], or generating plausible walking and sitting animations [31, 83, 77, 75, 76, 79, 103, 104, 99, 36, 107, 27]. This enables more natural modeling of human reactions to their environment; however, the generated interactions remain limited due to the assumption of a static scene environment, resulting in a focus on walking or sitting movements.
Recent methods have also focused on more fine-trained interactions by generating human motion given a single static object [67, 69, 98, 82, 42, 97, 41]. While these methods only focus on human motion generation for a static object, InterDiff [84] jointly forecasts both human and object motion sequences given an initial sequence observation. Our approach also models both human and object motion, but we formulate a flexible text-conditioned generative model for dynamic human and object motion, modeling the interdependency between human, object, and contact to synthesize more realistic interactions under various application settings.
Contact Prediction for Human-Object Interactions.
While there is a large corpus of related work for human motion prediction, only few works focus on object motion generation [19, 51, 59, 110]. Notably, these methods predict object movement in isolation, making interactions limited, as they typically involve interdependency with human motion.
@@ 4.1 Probabilistic Denoising Diffusion
Human-Object Interactions
To model human-object interactions with diffusion, we employ our neural network formulation $\mathcal{G}$. $\mathcal{G}$ operates on the noised vector of concatenated human, object, and contact representations, together with the current time step $t$, and a condition consisting of object point cloud $G$, encoded by an encoder $E_{G}$, and text information $T$, encoded by encoder $E_{T}$. Formally,
$\mathbf{\tilde{z}}=\mathcal{G}(\mathbf{z}_{t},t,E_{G}(G)\oplus E_{T}(T))$ (3)
@@
Object transformations $o_{i}$ are represented as global translation and rotation using continuous 6D rotation representation [109]. In contrast to prior work [91, 17, 71, 78, 94, 97, 41] which focused on representing human motion in a simplified manner as a collection of $J$ human joints, disregarding both identity-specific and pose-specific body shape, we model physically plausible human-object contacts between body surface and geometry. Thus, we represent the human body $h_{i}$ in SMPL [44] parameters: $h_{i}=\{h_{i}^{p},h_{i}^{b},h_{i}^{r},h_{i}^{t}\}$ with pose parameters $h_{i}^{p}\in\mathbb{R}^{63}$, shape parameters $h_{i}^{b}\in\mathbb{R}^{10}$, as well as global rotation $h_{i}^{r}\in\mathbb{R}^{3}$ and translation $h_{i}^{t}\in\mathbb{R}^{3}$. These body parameters can then be converted back into a valid human body surface mesh in a differentiable manner, using the SMPL [44] model. This allows us to reason about the contact between human body surface and object geometry. We represent contact $c_{i}$ on the human body as the distance between a set of $M=128$ uniformly distributed motion markers on the body surface to the closest point of the object geometry, for each marker. Specifically, we represent contact for frame $x_{i}$ and $j$-th contact marker ($j\in\{0..M\}$) $c_{i}^{j}$ as its distance from the human body surface to the closest point on the same frame’s object surface.
@@ 4.2 Human-Object-Contact Cross-Attention
We jointly predict human body sequences $\{h_{i}\}$, object transformations $\{o_{i}\}$, and corresponding contact distances $\{c_{i}\}$ in our diffusion approach. We employ a U-Net backbone for diffusion across these outputs, with separate residual blocks for human, object, and contact representations, building modality-specific latent feature representations. As we aim to model the inter-dependency across human, object, and contact, we introduce custom human-object-contact cross-attention modules after every residual block where each modality attends to the other two.
@@
with query $\boldsymbol{Q}=h_{i}$, and key and value $\boldsymbol{K}=\boldsymbol{V}=o_{i}\odot c_{i}$ ($\odot$ denotes concatenation). Applying this similarly to $o_{i}$ and $c_{i}$ yields the final features after each cross-attention module.
@@ 4.3 Contact-Based Object Transform Weighting
As visualized in Fig. 3, object motion is naturally most influenced by parts of the human body in very close contact to the object (as they are often the cause of the motion), and less impacted (if at all) by body parts further away. For instance, if a person moves an object with their hands, the object follows the hands but not necessarily other body parts (e.g., body and feet may remain static or walk in a different direction). Thus, instead of directly generating one object motion hypothesis $o_{i}$ alongside the corresponding human motion $h_{i}$, we couple $o_{i}$ to the $M$ body contact points $j\in\{1..M\}$ and their predicted distances $\{c_{i}^{j}\}$ between human body surface and object geometry.
@@ 4.4 Loss Formulation
During training, the input is a noised vector $\mathbf{z}$, containing $F$ frames $\{\mathbf{x}_{i}\}$, each a concatenation of human body representation $h_{i}$, object transformation $o_{i}$, and contact parameters $c_{i}$. As condition $\mathbf{C}$, we use additionally input encoded object geometry $G$ and text description $T$. The training process is then supervised with the ground-truth sequence containing $\hat{h_{i}},\hat{o_{i}},\hat{c_{i}}$, minimizing a common objective:
@@ 5 Interaction Generation
Specifically, we use our trained model to reverse the forward diffusion process of Eq. 2: Starting with noised sample $\mathbf{z}_{T}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$, we iteratively use our trained network model $\mathcal{G}$ to estimate cleaned sample $\mathbf{z}_{0}$:
@@ 5.1 Contact-Based Diffusion Guidance
While our joint human-object-contact training already leads to some plausible motions, generated sequences are not explicitly constrained to respect contact estimates during inference, which can lead to inconsistent contact between human and object motion (e.g., floating objects). Thus, we introduce a contact-based guidance during inference to refine predictions, using a cost function $G(\mathbf{(}x)_{t})$ which takes as input the denoised human, object, and contact predictions $\mathbf{z}_{t}=[h_{t},o_{t},c_{t}]$ at diffusion step $t$. Based on the difference between predicted and actual contact distances for each contact point, we then calculate the gradient $\nabla_{\mathbf{z}_{t}}G(\mathbf{z}_{t})$.
@@ 6 Results
Table 1: Quantitative comparison with state-of-the-art approaches, MDM [71] and InterDiff [84]. Human Only results are evaluated only on the human pose sequence, and motion-cond. denotes predictions additionally conditioned on past observations of both human and object behavior. For metrics with $\rightarrow$, results closer to the real distribution are better. Our approach outperforms these baselines in all three settings, indicating a strong learned correlation between human and object motion.
@@ 6.1 Experimental Setup
Datasets
We conduct our experiments on two datasets containing interactions between whole-body 3D humans and corresponding objects. CHAIRS [35] captures 46 subjects as their SMPL-X [53] bodies interacting with 81 different types of chairs and sofas. We extract sequences in which both human and object are in motion, yielding $\approx 1300$ HOI sequences, each labeled with a text description. We use a random 80/10/10 split along object classes, ensuring that test objects are not seen during training. BEHAVE [8] captures 8 participants as their SMPL-H [62] parameters alongside 20 different objects. This yields $\approx 520$ sequences with corresponding text descriptions. We use their original train/test split. We sample both datasets at 20 frames per second, and generate 32 frames for CHAIRS and 64 for BEHAVE, leading to generated motion that lasts up to 3 seconds.
Implementation Details We train our model with batch size 64 for 600k steps ($\approx$24 hours), after which we choose the checkpoint that minimized validation FID, following [84]. Our attention uses 4 heads and a latent dimension of 256. Input text is encoded using a frozen CLIP-ViT-B/32 model. For classifier-free guidance during inference time, we use a guidance scale of 2.5, which empirically provides a good trade-off between diversity and fidelity. For our inference-time contact-based guidance, we use scale $s=100.0$.
@@ 6.2 Evaluation Metrics
We measure realism and diversity of combined human and object motion, alongside closeness to the text description, following established practices [71, 25, 24]. We first train a joint human-object motion feature extractor and separate text feature extractor using a contrastive loss to produce geometrically close feature vectors for matched text-motion pairs, and vice versa. These encoders are then used for the following metrics:
@@
Perceptual User Study. The exact perceptual quality of human-object interactions is difficult to capture with any single metric; thus, we conducted a user study with 32 participants to evaluate our method in comparison to baseline approaches. Participants are shown side-by side views of sequences with the same geometry and text conditioning, and asked to choose 1) Which one follows the given text better and 2) Which one looks more realistic overall.
@@ 6.3 Comparison to Baselines
As our method is the first to enable generationg human and object motion from text, there are no baselines available for direct comparison. InterDiff [84] is closest to our approach, performing forecasting from observed human and object motion as input and predicting a plausible continuation. In Tab. 1, we compare to ours first in their setting, using observed motion as condition (motion-cond.), for a fair comparison. Additionally, we modify their approach by replacing observed motion encoders with our text encoder, allowing for a comparison in our setting (text-cond.). We also compare with MDM [71], a state-of-the-art method for human-only sequence generation from text, both in their original setting, only predicting human sequences, and extending theirs to also generate object sequences, by adding additional tokens and geometry conditioning to their transformer encoder formulation. For more details of baseline setup, we refer to the appendix. We evaluate the quality of generated human-object interactions as well as human-only generation, only evaluating the human sequence for our method, as compared to the generated sequences of MDM.
Both Tab. 1 and the user study in Tab. 5 show that our approach is able to generate more realistic and physically plausible human-object interaction sequences than baselines. In Fig. 4, we see that our approach synthesizes more meaningful human-object interaction with respect to contact and mitigating independent object floating.
@@ 6.4 Ablation Studies
Figure 6: Visualization of ablation of our method design: Generation, weighting, and inference-time guidance work together to enable realistic interactions in our method, resolving artifacts such as object floating.
Cross-attention enables learning human-object interdependencies.
Tab. 2 shows that our human-object-contact cross-attention (Sec. 4.2) significantly improves performance by effectively sharing information between human, contact, and object sequence modalities. In Fig. 6, we see this encourages realistic contact between human and object.
Contact prediction improves HOI generation performance.
Predicting contact (Sec. 5) is crucial to generating more realistic human-object sequences, resulting in more realistic interactions between human and object (Fig. 6), and improved fidelity (Tab. 2). Notably, learning contact jointly with human and object motion improves overall quality, compared to a separately trained contact model used for inference guidance (“Separate contact pred.”, Tab. 2).
Contact-based object transformation weighting improves generation performance.
Weighting predicted object motion hypotheses with predicted contact (Sec. 4.3) improves HOI generation over naive object sequence prediction, both quantitatively in Tab. 2 (“No contact weighting”) and visually as realistic human-object interactions in Fig. 6.
Contact-based guidance during inference helps produce physically plausible interactions.
As visualized in Fig. 6 and evaluated in Tab. 2, using custom guidance based on predicted contacts leads to a higher degree of fidelity and physical plausibility.

### v2

@@ Abstract
We propose CG-HOI, the first method to address the task of generating dynamic 3D human-object interactions (HOIs) from text. We model the motion of both human and object in an interdependent fashion, as semantically rich human motion rarely happens in isolation without any interactions. Our key insight is that explicitly modeling contact between the human body surface and object geometry can be used as strong proxy guidance, both during training and inference. Using this guidance to bridge human and object motion enables generating more realistic and physically plausible interaction sequences, where the human body and corresponding object move in a coherent manner. Our method first learns to model human motion, object motion, and contact in a joint diffusion process, inter-correlated through cross-attention. We then leverage this learned contact for guidance during inference to synthesize realistic and coherent HOIs. Extensive evaluation shows that our joint contact-based human-object interaction approach generates realistic and physically plausible sequences, and we show two applications highlighting the capabilities of our method. Conditioned on a given object trajectory, we can generate the corresponding human motion without re-training, demonstrating strong human-object interdependency learning. Our approach is also flexible, and can be applied to static real-world 3D scene scans.
@@ 1 Introduction
We present CG-HOI, the first approach to address the task of generating realistic 3D human-object interactions from text descriptions, by jointly predicting a sequence of 3D human body motion along with the object motion. Key to our approach is to not only model human and object motion, but to also explicitly model contact as a bridge between human and object. In particular, we model contact by predicting contact distances from the human body surface to the closest point on the surface of the object being interacted with. This explicit modeling of contact helps to encourage human and object motion to be semantically coherent, as well as to provide a constraint indicating physical plausibility (e.g., discouraging objects to float without support).
@@ 2 Related Work
Figure 2: Method Overview. Given a text description and object geometry, CG-HOI produces a human-object interaction (HOI) sequence, modeling both human and object motion. To produce realistic HOIs, we additionally model contact to bridge the interdependent motions. Our method jointly generates all three during training (left), using a U-Net-based diffusion with cross-attention across human, object, and contact. During inference (right), we drive synthesis under guidance of estimated contact to sample more physically plausible interactions.
3D Human Motion Generation. Generating sequences of 3D humans in motion is a task which evolved noticeably over the last few years. Traditionally, many methods used recurrent approaches [21, 2, 15, 23, 33, 52] and, improving both fidelity and predicted sequence length, graph- and attention-based frameworks [47, 48, 70]. Notably, generation can either happen deterministically, predicting one likely future human pose sequence [21, 52, 47, 48, 19], or stochastically, thereby also modelling the uncertainty inherent to future human motion [4, 7, 49, 90, 95, 18, 10, 89].
@@
3D Human Motion in Scenes. As human motion typically occurs not in isolation but in the context of an object or surrounding environment, various methods have explored learning plausible placement of humans into scenes, both physically and semantically, [100, 27, 104, 29, 30, 87], forecasting future motion given context [12, 50], or generating plausible walking and sitting animations [32, 86, 80, 78, 79, 82, 107, 108, 103, 37, 111, 28]. This enables more natural modeling of human reactions to their environment; however, the generated interactions remain limited due to the assumption of a static scene environment, resulting in a focus on walking or sitting movements.
Recent methods have also focused on more fine-grained interactions by generating human motion given a single static object [69, 71, 102, 85, 43, 101, 42]. While these methods only focus on human motion generation for a static object, [44] generates human motion conditioned on object motion and [88, 77] generate full human-object interaction sequences directly from an initial sequence observation. Our approach also models both human and object motion, but we formulate a flexible text-conditioned generative model for dynamic human and object motion, modeling the interdependency between human, object, and contact to synthesize more realistic interactions under various application settings.
Contact Prediction for Human-Object Interactions. While there is a large corpus of related work for human motion prediction, only few works focus on object motion generation [20, 53, 61, 114]. Notably, these methods predict object movement in isolation, making interactions limited, as they typically involve interdependency with human motion.
@@ 4.1 Probabilistic Denoising Diffusion
Human-Object Interactions To model human-object interactions with diffusion, we employ our neural network formulation $\mathcal{F}$. $\mathcal{F}$ operates on the noised vector of concatenated human, object, and contact representations, together with the current time step $t$, and a condition consisting of object point cloud $G$, encoded by an encoder $E_{G}$, and text information $T$, encoded by encoder $E_{T}$. Formally,
$\mathbf{\tilde{z}}=\mathcal{F}(\mathbf{z}_{t},t,E_{G}(G)\oplus E_{T}(T))$ (3)
@@
Object transformations $o_{i}$ are represented as global translation and rotation using continuous 6D rotation representation [113]. In contrast to prior work [95, 18, 73, 81, 98, 101, 42] which focused on representing human motion in a simplified manner as a collection of $J$ human joints, disregarding both identity-specific and pose-specific body shape, we model physically plausible human-object contacts between body surface and geometry. Thus, we represent the human body $h_{i}$ in SMPL [46] parameters: $h_{i}=\{h_{i}^{p},h_{i}^{b},h_{i}^{r},h_{i}^{t}\}$ with pose parameters $h_{i}^{p}\in\mathbb{R}^{63}$, shape parameters $h_{i}^{b}\in\mathbb{R}^{10}$, as well as global rotation $h_{i}^{r}\in\mathbb{R}^{3}$ and translation $h_{i}^{t}\in\mathbb{R}^{3}$. These body parameters can then be converted back into a valid human body surface mesh in a differentiable manner using the SMPL [46] model. This allows us to reason about the contact between human body surface and object geometry. We represent contact $c_{i}$ on the human body as the distance between a set of $M=128$ uniformly distributed motion markers on the body surface to the closest point of the object geometry, for each marker. Specifically, we represent contact for frame $\mathbf{x}_{i}$ and $j$-th contact marker ($j\in\{0..M-1\}$) $c_{i}^{j}$ as its distance from the human body surface to the closest point on the same frame’s object geometry surface.
@@ 4.2 Human-Object-Contact Cross-Attention
We jointly predict human body sequences $H=\{h_{i}\}$, object transformations $O=\{o_{i}\}$, and corresponding contact distances $C=\{c_{i}\}$ in our diffusion approach. We employ a U-Net backbone for diffusion across these outputs, with separate residual blocks for human, object, and contact representations, building modality-specific latent feature representations. As we aim to model the inter-dependency across human, object, and contact, we introduce custom human-object-contact cross-attention modules after every residual block where each modality attends to the other two.
@@
with query $\boldsymbol{Q}=H$, and key and value $\boldsymbol{K}=\boldsymbol{V}=O\odot C$ ($\odot$ denotes concatenation), i.e. $\boldsymbol{Q}\in\mathbb{R}^{F\times d}$ and $\boldsymbol{K},\boldsymbol{V}\in\mathbb{R}^{2F\times d}$. As in [76], $d$ denotes the dimensionality of query and key. Applying this similarly to $O$ and $C$ yields the final features after each cross-attention module.
@@ 4.3 Contact-Based Object Transform Weighting
As visualized in Fig. 3, object motion is naturally most influenced by parts of the human body in very close contact to the object (as they are often the cause of that motion), and less impacted (if at all) by body parts further away. For instance, if a person moves an object with their hands, the object follows the hands but not necessarily other body parts (e.g., body and feet may remain static or walk in a different direction). Thus, instead of directly generating one object motion hypothesis $o_{i}$ alongside the corresponding human motion $h_{i}$, we couple $o_{i}$ to the $M$ body contact points $j\in\{0..M-1\}$ and their predicted distances $\{c_{i}^{j}\}$ between human body surface and object geometry.
@@ 4.4 Loss Formulation
During training, the input is a noised vector $\mathbf{z}$, containing $F$ frames $\{\mathbf{x}_{i}\}$, each a concatenation of human body representation $h_{i}$, object transformation $o_{i}$, and contact parameters $c_{i}$. As condition $\mathbf{C}$, we additionally input encoded object geometry $G$ and text description $T$. The training process is then supervised with the ground-truth sequence containing $\hat{h_{i}},\hat{o_{i}},\hat{c_{i}}$, minimizing a common objective:
@@ 5 Interaction Generation
Specifically, we use our trained model to reverse the forward diffusion process of Eq. 2: Starting with noised sample $\mathbf{z}_{T}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$, we iteratively use our trained network model $\mathcal{F}$ to estimate cleaned sample $\mathbf{z}_{0}$:
@@ 5.1 Contact-Based Diffusion Guidance
While our joint human-object-contact training already leads to plausible motions, generated sequences are not explicitly constrained to respect contact estimates during inference, which can lead to inconsistent contact between human and object motion (e.g., floating objects). Thus, we introduce a contact-based guidance during inference to refine predictions, using a cost function $\mathcal{G}(\mathbf{z}_{t})=||c_{t}-\overline{c}_{t}||_{2}^{2}$ which takes as input the denoised human, object, and contact predictions $\mathbf{z}_{t}=[h_{t},o_{t},c_{t}]$ at diffusion step $t$ and compares predicted $c_{t}$ and actual contact distances $\overline{c}_{t}$ for each contact point. Based on this, we then calculate the gradient $\nabla_{\mathbf{z}_{t}}\mathcal{G}(\mathbf{z}_{t})$.
@@ 6 Results
Table 1: Quantitative comparison with state-of-the-art approaches MDM [73] and InterDiff [88]. Human Only results are evaluated only on the human pose sequence, and motion-cond. denotes predictions additionally conditioned on past observations of both human and object behavior. For metrics with $\rightarrow$, results closer to the real distribution are better. Our approach outperforms these baselines in all three settings, indicating a strong learned correlation between human and object motion.
@@ 6.1 Experimental Setup
Datasets We conduct our experiments on two datasets containing interactions between whole-body 3D humans and corresponding objects. CHAIRS [36] captures 46 subjects as their SMPL-X [55] bodies interacting with 81 different types of chairs and sofas. We extract sequences in which both human and object are in motion, yielding $\approx 1300$ HOI sequences, each labeled with a text description. We use a random 80/10/10 split along object classes, ensuring that test objects are not seen during training. BEHAVE [9] captures 8 participants as their SMPL-H [64] parameters alongside 20 different objects. This yields $\approx 520$ sequences with corresponding text descriptions. We use their original train/test split. We sample both datasets at 20 frames per second, and generate 32 frames for CHAIRS and 64 for BEHAVE, leading to generated motion that lasts up to 3 seconds.
Implementation Details We train our model with batch size 64 for 600k steps ($\approx$24 hours), after which we choose the checkpoint that minimizes validation FID, following [88]. Our attention uses 4 heads and a latent dimension of 256. Input text is encoded using a frozen CLIP-ViT-B/32 model. For classifier-free guidance during inference time, we use a guidance scale of 2.5, which empirically provides a good trade-off between diversity and fidelity. For our inference-time contact-based guidance, we use scale $s=100.0$.
@@ 6.2 Evaluation Metrics
We measure realism and diversity of combined human and object motion, alongside closeness to the text description, following established practices [73, 26, 25]. We first train a joint human-object motion feature extractor and separate text feature extractor using a contrastive loss to produce geometrically close feature vectors $\{v_{i}\}$ for matched text-motion pairs, and report the following metrics:
@@
Perceptual User Study. The exact perceptual quality of human-object interactions is difficult to capture with any single metric; thus, we additionally conduct a user study with 32 participants to evaluate our method in comparison to baseline approaches. Participants are shown 10 baseline vs. ours pairs each in side-by side views of sequences with the same geometry and text conditioning, and asked to choose 1) Which one follows the given text better and 2) Which one looks more realistic overall.
@@ 6.3 Comparison to Baselines
As our method is the first to enable generating human and object motion from text, there are no baselines available for direct comparison. InterDiff [88] is closest to our approach, performing forecasting from observed human and object motion as input and predicting a plausible continuation. In Tab. 1, we compare to ours first in their setting, using observed motion as condition (motion-cond.), for a fair comparison. Additionally, we modify their approach by replacing observed motion encoders with our text encoder, allowing for a comparison in our setting (text-cond.). We also compare with MDM [73], a state-of-the-art method for human-only sequence generation from text, both in their original setting, only predicting human sequences, and extending theirs to also generate object sequences, by adding additional tokens and geometry conditioning to their transformer encoder formulation. For more details of baseline setup, we refer to the appendix. We evaluate the quality of generated human-object interactions as well as human-only generation, only evaluating the human sequence for our method, as compared to the generated sequences of MDM.
Both Tab. 1 and the user study in Fig. 5 show that our approach is able to generate more realistic and physically plausible human-object interaction sequences than baselines. In Fig. 4, we see that our approach synthesizes more meaningful human-object interaction with respect to contact and mitigating independent object floating.
@@ 6.4 Ablation Studies
Figure 6: Visualization of ablations of our method design: Generation, weighting, and inference-time guidance work together to enable realistic interactions in our method, resolving artifacts such as object floating.
Cross-attention enables learning human-object interdependencies. Tab. 2 shows that our human-object-contact cross-attention (Sec. 4.2) significantly improves performance by effectively sharing information between human, contact, and object sequence modalities. In Fig. 6, we see this encourages realistic contact between human and object.
Contact prediction improves HOI generation performance. Predicting contact (Sec. 5) is crucial to generating more realistic human-object sequences, resulting in more realistic interactions between human and object (Fig. 6), and improved fidelity (Tab. 2). Notably, learning contact jointly with human and object motion improves overall quality, compared to a separately trained contact model used for inference guidance (“Separate contact pred.”, Tab. 2).
Contact-based object transformation weighting improves generation performance. Weighting predicted object motion hypotheses with predicted contact (Sec. 4.3) improves HOI generation over naive object sequence prediction, both quantitatively in Tab. 2 (“No contact weighting”) and visually as realistic human-object interactions in Fig. 6.
Contact-based guidance during inference helps produce physically plausible interactions. As shown in Fig. 6 and Tab. 2, using our guidance based on predicted contacts leads to a higher degree of fidelity and physical plausibility.
6.5 Applications
Human motion generation given object trajectory. Our approach can be directly applied to conditionally generate human sequences given object sequences as condition, as shown in Fig. 7. As our model learns a strong correspondence between object and human motion, facilitated by contact distance predictions, we are able to condition without any additional training.